I can not figure this out please help me. Thanks so much Place the following set of numbers in ascending order.

,5/9,1/8,,2/7

Answer by edjones(8007)

Easiest is to turn them into decimals and then arrange them, lowest to highest.
-2/3=-.67
-5/9=-.56
1/8=.125
2/7=.29
5/9=.56
Ed